OpenStudy (brooke..help00):

WILL FAN AND MEDAL! Solve for x: 2 over 3 (x − 4) = 2x. 2 −2 −8 −4

OpenStudy (brooke..help00):

@Will.H

OpenStudy (brooke..help00):

@Wholock

OpenStudy (1davey29):

(2/3)(x-4)=2x, solve for x, distribute 2/3, (2x-8)/3=2x, split the fraction, ((2x)/3)-(8/3)=2x, 8/3=2x-((2x)/3), 8/3=((6x)/3)-((2x)/3), 8/3=((4x)/3), 8=4x, x=2

OpenStudy (brooke..help00):

That doesn't make since can you split it up?

OpenStudy (otherworldly):

2/3 = 0.6666......

OpenStudy (vjt):

is this how your equation looks? \[2\div3(x-4)=2x\]

OpenStudy (otherworldly):

thats the equation yup

OpenStudy (1davey29):

If you have one term in the denominator and two or more in the numerator separated by a plus or minus, you can split the fraction by keeping the denominator for both and separating the numerators into their own fractions with the sign that was between them between the fractions.

jhonyy9 (jhonyy9):

2 ---(x-4) = 2x multiply both sides by 3 and will get 3 2(x-4)=6x do you can continue it from here ?

jhonyy9 (jhonyy9):

hi i like help you please collaborate

jhonyy9 (jhonyy9):

ok. ?

jhonyy9 (jhonyy9):

2(x-4)=6x next step divide both sides by 2 so what will get ?

OpenStudy (brooke..help00):

2x -8 = 6x -- -- 2 2 x-8=3x? Confused again:/

jhonyy9 (jhonyy9):

no there are 2(x-4)=6x and so divide both sides by 2

OpenStudy (wholock):

Okay brooke, stop right there, DO NOT DIVIDE 2

OpenStudy (wholock):

Use the distributive property first 2x - 8 = 6x From there, subtract 2x from both sides to get 4x = -8 Divide both sides by 4 to get x = -2

OpenStudy (wholock):

And there you go @brooke..help00 . got it?

OpenStudy (brooke..help00):

Yes! Thank you I have a lot of question... an you help?

OpenStudy (1davey29):

You guys are forgetting it was (2/3)(x-4)=2x, not (2(x-4))/3=2x, so you can't divide the 3 first.